A family has two cars. The first car has a fuel efficiency of 40 miles per gallon of gas and the second has a fuel efficiency of 15 miles per gallon of gas. during one particular week, the two cars went a combined total of 1250 miles, for a total gas consumption of 50 gallons. how many gallons were consumed by each of the two cars that week?
Answer by Boreal(15235)

first car consumed x gallons
second car consumed the rest, 50-x
miles were 40x+15(50-x)=1250
40x+750-15x=1250
25x=500
x=20 gallons for first car and 800 miles
50-x=30 gallons for second car and 450 miles
They add to 1250 miles